OpenStudy (anonymous):

Which of the following is an irrational number? a 9 squared b 84 squared c 1.7 d 64 squared

jimthompson5910 (jim_thompson5910):

I'm assuming you meant to say \[\Large \sqrt{9}\] for choice A and similar for B & D

OpenStudy (anonymous):

you'd be correct then

jimthompson5910 (jim_thompson5910):

if you take the square root of a positive whole number and you get a whole number, then that original number is rational. Otherwise, it's irrational

jimthompson5910 (jim_thompson5910):

example \[\Large \sqrt{16} = 4\] so that is rational

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{23} \approx 4.79583\] (the decimals go on forever with no pattern) so that is irrational

OpenStudy (anonymous):

so how would i solve this problem for example how should i divide it

jimthompson5910 (jim_thompson5910):

take the square root of the stuff under the square root symbol and use the rule I mentioned above

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then 84 is the correct answer then right?

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

yes because \[\Large \sqrt{84} \approx 9.16515\] while the other roots are whole numbers

jimthompson5910 (jim_thompson5910):

the 1.7 is not irrational because you can say 1.7 = 17/10